OpenStudy (foreverandalways13):

a ski chalet offers 2 weekend packages. one package provides 2 nights and 3 meals for$280. the other provides 2 nights and 5 meals for $320. write and solve a system of equations to find the cost per night and per meal. HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

2n+3m=280; 2n+5m=320. Substract both equations: 2m=40, then m=20. Replace it and 2n+60=280, 2n=220, n=110. Cost per night=$110; cost per meal=$20.

OpenStudy (foreverandalways13):

THANK YOU!!!

OpenStudy (anonymous):

u'r welcome. Greetings from La Paz, Bolivia

OpenStudy (foreverandalways13):

so cool... haha im from California usa!

OpenStudy (foreverandalways13):

*USA